Find the transformation of the equation 12x2 −2y2 +z2 = 2xy if the origin is kept fixed and the axes are rotated in such a way that the direction ratios of the new axes are 1,−3,0; 3,1,0; 0,0,1.


1
Expert's answer
2021-06-03T03:54:36-0400

The normalized axes are

"e'_1=\frac{1}{\sqrt{10}}(1,-3,0),e'_2=\frac{1}{\sqrt{10}}(3,1,0),e'_3=(0,0,1)"

or

"\begin{pmatrix}
 e'_1 \\
 e'_2\\
e'_3
\end{pmatrix}=\begin{pmatrix}
 \frac{1}{\sqrt{10}} & \frac{-3}{\sqrt{10}}&0 \\
 \frac{3}{\sqrt{10}} & \frac{1}{\sqrt{10}}&0\\
0&0&1
\end{pmatrix}\begin{pmatrix}
 e_1 \\
 e_2\\
e_3
\end{pmatrix}"

- the basis in the rotated frame. The transformation of coordinates is given by

"\begin{pmatrix}
 x' \\
 y'\\
z'
\end{pmatrix}=\begin{pmatrix}
 \frac{1}{\sqrt{10}} & \frac{-3}{\sqrt{10}}&0 \\
 \frac{3}{\sqrt{10}} & \frac{1}{\sqrt{10}}&0\\
0&0&1
\end{pmatrix}^{-1}\begin{pmatrix}
 x \\
 y\\
z
\end{pmatrix}"

So:

"\begin{pmatrix}
 x\\
 y\\
z
\end{pmatrix}=\begin{pmatrix}
 \frac{1}{\sqrt{10}} & \frac{-3}{\sqrt{10}}&0 \\
 \frac{3}{\sqrt{10}} & \frac{1}{\sqrt{10}}&0\\
0&0&1
\end{pmatrix}\begin{pmatrix}
 x' \\
 y'\\
z'
\end{pmatrix}"

"x = \begin{cases}
 x=\frac{x'}{\sqrt{10}}-\frac{3y'}{\sqrt{10}} \\
 y=\frac{3x'}{\sqrt{10}}+\frac{y'}{\sqrt{10}}\\
z=z'
\end{cases}"

"12x^2-2y^2+z^2=2xy"


"\frac{12(x'-3y')^2}{10}-\frac{2(3x'-y')^2}{10}+z'^2=\frac{2(x'-3y')(3x'-y')}{10}"


"12(x'^2-6x'y'+9'^2)-2(9x'^2+6x'y'+y'^2)+10z'^2=2(3x'^2-8x'y'-3y'^2)"


"-12x'^2-68x'y'+112y'^2+10z'^2=0"


Answer:


"5z'^2+56y'^2-6x'^2=34x'y'"